# Finite energy well-posedness for nonlinear Schrödinger equations with non-vanishing conditions at infinity

### v3

@@ 2.2. The Hamiltonian for wave-functions in the energy space
is introduced. It follows that $(\mathbb{E},d_{\mathbb{E}}^{\prime})$ is connected. Relying on the decomposition of elements of $\mathbb{E}$ provided by [24, Theorem 1.8], one can show that $(\mathbb{E},d_{\mathbb{E}}^{\prime})$ is separable. If one only requires continuity of the solution map with respect to this weakened topology, the proof of Proposition 3.2 can be simplified. This metric has widely been used in the study of the stability of special solutions for $d=1$. We refer to [48], where the authors introduce new energy spaces for (1.6) and $d=1$ in order to tackle global well-posedness in the energy space at $H^{s}$-regularity.
@@
$\||\psi_{1}|-|\psi_{2}|\|_{L^{2}({\mathbf{R}}^{d})}\leq C\left(\||\psi_{1}|^{2}-|\psi|^{2}\|_{L^{2}({\mathbf{R}}^{d})}+\|\nabla\psi_{1}-\nabla\psi_{2}\|_{L^{2}({\mathbf{R}}^{d})}\right).$
@@
For $d=3$, one proceeds similarly exploiting the decomposition $\psi_{i}=c_{i}+v_{i}$, $v_{i}\in\mathcal{F}_{c}({\mathbf{R}}^{3})$ and Proposition 2.2. It holds
@@
$\||\psi_{1}|^{2}(1-\chi(\psi_{1}))-|\psi_{2}|^{2}(1-\chi(\psi_{2}))\|_{L^{2}({\mathbf{R}}^{d})}\\ \leq C\left(1+\sqrt{\mathcal{E}(\psi_{1})}+\sqrt{\mathcal{E}(\psi_{2})}+\|\psi_{1,q}\|_{L^{4}({\mathbf{R}}^{d})}+\|\psi_{2,q}\|_{L^{4}({\mathbf{R}}^{d})}\right)\|\psi_{1,q}-\psi_{2,q}\|_{L^{4}({\mathbf{R}}^{d})}\\ \leq 2C\left(1+\sqrt{\mathcal{E}(\psi_{1})}+\sqrt{\mathcal{E}(\psi_{2})}\right)\|\psi_{1,q}-\psi_{2,q}\|_{L^{4}({\mathbf{R}}^{d})}.$
@@
(2.15) $|\psi|^{4}\sqrt{1-\chi(\psi)}\leq C\left|\psi_{q}\right|^{4},$
with $\psi_{q}$ defined in (2.2) which is only valid provided $(1-\chi(\psi))>\theta$ for some small $\theta>0$. However, this is harmless as
[…]
@@
$\|\psi_{1,q}-\psi_{2,q}\|_{L^{4}({\mathbf{R}}^{d})}\leq C\left(\sqrt{\mathcal{E}(\psi_{1})}+\sqrt{\mathcal{E}(\psi_{2})}\right)\|\psi_{1}-\psi_{2}\|_{X^{1}+H^{1}({\mathbf{R}}^{d})}$
@@
It remains to show (2.14). The respective property is known for $d_{\mathbb{E}_{\mathrm{GL}}}$, see [23, Lemma 2], and hence follows from the equivalence of metrics. However, we provide a proof to track constants explicitly. Note that
@@
Next, we provide a sufficient condition on $f$ under which the space of functions with finite Hamiltonian energy is included in $\mathbb{E}$. To that end, we require Assumption 1.5 to be satisfied. From $F(1)=F^{\prime}(1)=f(1)=0$ and Taylor expansion it follows
(2.16) $F(r)\simeq\frac{1}{2}f^{\prime}(1)(r-1)^{2}$
in a small neighborhood of $1$. Hence, there exists $\delta>0$ such that for all $r\in(1-\delta,1+\delta)$ there exists $C_{1},C_{2}>0$ such that
provided that $||\psi|^{2}-1|<\delta$. The nonlinear potential energy density $F$ is locally convex in a neighborhood of $1$. It was shown in [17, Lemma 4.8] that requiring in addition that $F\geq 0$ and hence the Hamiltonian energy is sign-definite, implies that $\mathbb{E}=\{\mathcal{H}(\psi)<\infty\}$. Note that the condition $F\geq 0$ is for instance satisfied for the pure power-type nonlinearities in (1.4).
@@
Let $d=2,3$ and Assumptions 1.5 be satisfied. If in addition $F\geq 0$, then
@@
Since $|\varphi(|\psi|)^{2}-1|\leq 4||\psi|-1|$, one has $\varphi(|\psi|)^{2}-1\in L^{2}({\mathbf{R}}^{d})$ if $\psi\in\mathbb{E}({\mathbf{R}}^{d})$. For the converse, see [17, Lemma 2.1]. We sketch the main idea. On the set where $|\psi(x)|\leq 2$, one has $\varphi(|\psi|)^{2}=|\psi|^{2}$ and hence the desired bound follows. Further, $\mathcal{L}^{d}(\{x\,:\,||\psi(x)|-1|>\frac{3}{2}\})<+\infty$ from the Chebychev inequality (2.8) if $\varphi(|\psi|)^{2}-1\in L^{2}({\mathbf{R}}^{d})$. By means of (2.11) for $d=2$ and Sobolev embedding for $d=3$ one concludes. Finally, there exists $C>0$ and an increasing function $m:{\mathbf{R}}_{+}\rightarrow{\mathbf{R}}_{+}$ with $\lim_{r\rightarrow 0}m(r)=0$ such that
@@
see [17, Corollary 4.3]. Second, we note that it suffices to establish inequality (2.18) for $\mathcal{E}$ replaced by $\mathcal{E}_{\mathrm{mGL}}$. In virtue of (2.17), it suffices to consider the region where $\{x:||\psi|-1|\geq\delta\}$. If $\inf F>0$ on $\{x:||\psi|-1|\geq\delta\}$, then it is clear that
@@
It follows that $\mathcal{E}(\psi)$ can be controlled in terms of $\mathcal{H}(\psi)$. More in general, provided that $F\geq 0$, it follows from [17, Lemma 4.8] that for all $\psi$ with $|\mathcal{H}(\psi)|<\infty$ there exist $C_{1}=C_{1}(\mathcal{H}(\psi))>0$ and $C_{2}=C_{2}(\mathcal{H}(\psi))>0$ such that
@@
System (1.1) is closely related to the QHD system with non-trivial far-field. In a reminiscent analysis, the regularity and integrability properties of its unknowns $(\rho,J)$ corresponding to the mass density $\rho=|\psi|^{2}$ and momentum density $J=\operatorname{Im}(\overline{\psi}\nabla\psi)$ are then captured in terms of Orlicz spaces, see [3] and [36, Chapter 2] as well as [2, 37] for the respective uniform bounds for solutions to the quantum Navier-Stokes equations, a viscous regularization of the QHD system.
@@ 2.3. Smooth approximation
The first statement is proven in [23, Lemma 6] by considering the convolution with a standard mollification kernel and the second statement follows from [22, Proposition 1.1.]. In [23, 22], the statements are given for $(\mathbb{E}_{\mathrm{GL}},d_{\mathbb{E}_{\mathrm{GL}}})$ being equivalent to $(\mathbb{E},d_{\mathbb{E}})$ by virtue of Lemma 2.6.
@@ 2.4. Action of the linear propagator on the energy space
The action of the linear Schrödinger group on the space $X^{k}({\mathbf{R}}^{d})+H^{k}({\mathbf{R}}^{d})$ is well-defined, see [23, Lemma 3] and also [24]. While the results in [23, 24] are stated for $(\mathbb{E}_{\mathrm{GL}},d_{\mathbb{E}_{\mathrm{GL}}})$, we state them $(\mathbb{E},d_{\mathbb{E}})$ which by Lemma 2.6 is equivalent.
@@
(2.20) $\|\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}f\|_{X^{k}+H^{k}}\leq C\left(1+t\right)^{\frac{1}{2}}\|f\|_{X^{k}+H^{k}},$
@@
The action of the linear Schrödinger group on the space $\mathbb{E}({\mathbf{R}}^{d})$ is described by [23, Proposition 2.3].
@@
Let $d=2,3$. For every $t\in{\mathbf{R}}$, the linear propagator $\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}$ maps $\mathbb{E}({\mathbf{R}}^{d})$ to itself and for every $\psi\in\mathbb{E}({\mathbf{R}}^{d})$ the map $t\in{\mathbf{R}}\mapsto\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}\in\mathbb{E}({\mathbf{R}}^{d})$ is continuous. Moreover, given $R>0$, $T>0$ there exists $C>0$ such that for every $\psi_{0}^{1},\psi_{0}^{2}\in\mathbb{E}({\mathbf{R}}^{d})$ with $\mathcal{E}(\psi_{0}^{1})\leq R$, $\mathcal{E}(\psi_{0}^{2})\leq R$ one has
@@
Further, given $R>0$, there exists $T(R)>0$ such that, for every $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{d})$ with $\mathcal{E}(\psi_{0})\leq R$, we have
@@ 2.5. Strichartz estimates
for any admissibile pair $(q_{1},r_{1})$. Further, in order to discuss the well-posedness theory for (1.1) in the energy space, we also work with the function space $S^{1}(I\times{\mathbf{R}}^{d})$ and $N^{1}(I\times{\mathbf{R}}^{d})$ defined by the norms
@@
Let $T>0$ and $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{d})$, then Lemma 2.14 states that for any admissible pair $(q,r)$ it holds
@@
In virtue of Lemma 2.11, one has $\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}-\psi_{0}\in C([0,T];H^{1}({\mathbf{R}}^{d}))$ and $\nabla\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}\in C([0,T];L^{2}({\mathbf{R}}^{d}))\cap S^{0}([0,T]\times{\mathbf{R}}^{d}))$.
@@ 2.6. The nonlinearity
(2.31) $\mathcal{N}_{2,\infty}=\mathcal{N}_{2}(\psi)\chi(2\psi),\quad\mathcal{N}_{2,q}(\psi)=\mathcal{N}_{2}(\psi)(1-\chi(2\psi))$
@@
$|\mathcal{N}_{2,\infty}(\psi)|$ $\leq C(1-\eta(|\psi|),\quad|\mathcal{N}_{2,q}(\psi)|\leq C|\psi|^{2\alpha+1}(1-\chi(\psi)).$
In the case of vanishing boundary conditions and infinity, the strategy developed in [39], see also [14, Chapter 4], relies on similar pointwise bounds on $\mathcal{N}$. However, here we need to consider additional cut-off functions $\eta$ isolating the behavior close to $1$ in view of the far-field and the related support properties. Note that (2.8) yields that the measure of $\operatorname{supp}(\mathcal{N}_{2}(\psi))$ is bounded by $\mathcal{E}(\psi)$. The quantity $\nabla\mathcal{N}$ can be rigorously defined by means of Nemicki operators, see [39, Appendix A] and also [40, 14]. It reads
@@
For general $\psi_{1},\psi_{2}\in\mathbb{E}({\mathbf{R}}^{d})$ one has $\psi_{1}-\psi_{2}\notin L^{p}({\mathbf{R}}^{d})$ for any $p\geq 1$, unless $\psi_{1},\psi_{2}$ belong to the same connected component of $\mathbb{E}({\mathbf{R}}^{d})$, see Remark 2.3 and 2.4 for $d=2,3$ respectively. This motivates the following estimates,
@@
$\left|\mathcal{N}_{2,\infty}(\psi_{1})-\mathcal{N}_{2,\infty}(\psi_{2})\right|$ $\leq C\left|\psi_{1}-\psi_{2}\right|,$
$\left|\mathcal{N}_{2,q}(\psi_{1})-\mathcal{N}_{2,q}(\psi_{2})\right|$ $\leq C\left(|\psi_{1}|^{2\alpha}+|\psi_{2}|^{2\alpha}\right)\left|\psi_{1}-\psi_{2}\right|.$
@@
(2.37) $G_{1}(\psi)=f(|\psi|^{2})+f^{\prime}(|\psi|^{2})|\psi|^{2},\quad G_{2}(\psi)=f^{\prime}(|\psi|^{2})\psi^{2}.$
@@
$G_{i,\infty}(\psi):=G_{i}(\psi)\chi(\psi),\quad G_{i,q}(\psi):=G_{i}(\psi)(1-\chi(\psi)),$
@@
(2.38) $G_{\infty}:=G_{1,\infty}+G_{2,\infty},\quad G_{q}:=G_{1,q}+G_{2,q}.$
@@
(2.39) $|G_{\infty}(\psi)|\leq C,\quad|G_{q}(\psi)|\leq C(1+|\psi_{q}|^{2\alpha})(1-\chi(\psi)).$

### v4

@@ 2.2. The Hamiltonian for wave-functions in the energy space
is introduced. It follows that $(\mathbb{E},d_{\mathbb{E}}^{\prime})$ is connected. Relying on the decomposition of elements of $\mathbb{E}$ provided by [27, Theorem 1.8], one can show that $(\mathbb{E},d_{\mathbb{E}}^{\prime})$ is separable. If one only requires continuity of the solution map with respect to this weakened topology, the proof of Proposition 3.2 can be simplified. This metric has widely been used in the study of the stability of special solutions for $d=1$. We refer to [51], where the authors introduce new energy spaces for (1.6) and $d=1$ in order to tackle global well-posedness in the energy space at $H^{s}$-regularity.
@@
$\||\psi_{1}|-|\psi_{2}|\|_{L^{2}({\mathbf{R}}^{d})}\leq C\left(\||\psi_{1}|^{2}-|\psi_{2}|^{2}\|_{L^{2}({\mathbf{R}}^{d})}+\|\nabla\psi_{1}-\nabla\psi_{2}\|_{L^{2}({\mathbf{R}}^{d})}\right).$
@@
For $d=3$, one proceeds similarly exploiting the decomposition $\psi_{i}=c_{i}+v_{i}$, $v_{i}\in\mathcal{F}_{c}({\mathbf{R}}^{3})$ and Proposition 2.2. We have
@@
$\||\psi_{1}|^{2}(1-\chi(\psi_{1}))-|\psi_{2}|^{2}(1-\chi(\psi_{2}))\|_{L^{2}({\mathbf{R}}^{d})}\\ \leq C\left(1+\sqrt{\mathcal{E}(\psi_{1})}+\sqrt{\mathcal{E}(\psi_{2})}+\|\psi_{1,\mathrm{int}}\|_{L^{4}({\mathbf{R}}^{d})}+\|\psi_{2,\mathrm{int}}\|_{L^{4}({\mathbf{R}}^{d})}\right)\|\psi_{1,\mathrm{int}}-\psi_{2,\mathrm{int}}\|_{L^{4}({\mathbf{R}}^{d})}\\ \leq 2C\left(1+\sqrt{\mathcal{E}(\psi_{1})}+\sqrt{\mathcal{E}(\psi_{2})}\right)\|\psi_{1,\mathrm{int}}-\psi_{2,\mathrm{int}}\|_{L^{4}({\mathbf{R}}^{d})}.$
@@
(2.16) $|\psi|^{4}\sqrt{1-\chi(\psi)}\leq C\left|\psi_{\mathrm{int}}\right|^{4},$
with $\psi_{\mathrm{int}}$ defined in (2.2) which is only valid provided $(1-\chi(\psi))>\theta$ for some small $\theta>0$. However, this is harmless as
$\mathcal{L}^{2}\left(\left\{x\in\operatorname{supp}(1-\chi(\psi))\,:\,0<1-\chi(\psi)\leq\theta\right\}\right)\leq\sqrt{\mathcal{E}(\psi)}$
@@
$\|\psi_{1,\mathrm{int}}-\psi_{2,\mathrm{int}}\|_{L^{4}({\mathbf{R}}^{d})}\leq C\left(\sqrt{\mathcal{E}(\psi_{1})}+\sqrt{\mathcal{E}(\psi_{2})}\right)\|\psi_{1}-\psi_{2}\|_{X^{1}+H^{1}({\mathbf{R}}^{d})}$
@@
It remains to show (2.15). The respective property is known for $d_{\mathbb{E}_{\mathrm{GL}}}$, see [26, Lemma 2], and hence follows from the equivalence of metrics. However, we provide a proof to track constants explicitly. Note that
@@
Next, we provide a sufficient condition on $f$ under which the space of functions with finite Hamiltonian energy is included in $\mathbb{E}$. To that end, we require Assumption 1.5 to be satisfied so that (2.12) holds. The nonlinear potential energy density $F$ is locally convex in a neighborhood of $1$. It was shown in [17, Lemma 4.8] that requiring in addition that $F\geq 0$ and hence the Hamiltonian energy is sign-definite, implies that $\mathbb{E}=\{\mathcal{H}(\psi)<\infty\}$. Note that the condition $F\geq 0$ is for instance satisfied for the pure power-type nonlinearities in (1.4).
@@
Let $d=2,3$ and suppose that Assumption 1.5 is satisfied. If in addition $F\geq 0$, then
@@
Since $|\varphi(|\psi|)^{2}-1|\leq 4||\psi|-1|$, one has $\varphi(|\psi|)^{2}-1\in L^{2}({\mathbf{R}}^{d})$ if $\psi\in\mathbb{E}({\mathbf{R}}^{d})$. For the converse, see [17, Lemma 2.1]. We sketch the main idea. On the set where $|\psi(x)|\leq 2$, one has $\varphi(|\psi|)^{2}=|\psi|^{2}$ and hence the desired bound follows. Further, $\mathcal{L}^{d}(\{x\,:\,||\psi(x)|-1|>\frac{3}{2}\})<+\infty$ from the Chebychev inequality (2.4) if $\varphi(|\psi|)^{2}-1\in L^{2}({\mathbf{R}}^{d})$. By means of (2.10) for $d=2$ and Sobolev embedding for $d=3$ one concludes. Finally, there exist $C>0$ and an increasing function $m:{\mathbf{R}}_{+}\rightarrow{\mathbf{R}}_{+}$ with $\lim_{r\rightarrow 0}m(r)=0$ such that
@@
see [17, Corollary 4.3]. Second, we note that it suffices to establish inequality (2.17) with $\mathcal{E}$ replaced by $\mathcal{E}_{\mathrm{mGL}}$. In virtue of (2.12), it suffices to consider the region where $\{x:||\psi|-1|\geq\delta\}$. If $\inf F>0$ on $\{x:||\psi|-1|\geq\delta\}$, then it is clear that
@@
It follows that $\mathcal{E}(\psi)$ can be controlled in terms of $\mathcal{H}(\psi)$. More generally, provided that $F\geq 0$, it follows from [17, Lemma 4.8] that for all $\psi$ with $|\mathcal{H}(\psi)|<\infty$ there exist $C_{1}=C_{1}(\mathcal{H}(\psi))>0$ and $C_{2}=C_{2}(\mathcal{H}(\psi))>0$ such that
@@
System (1.1) is closely related to the QHD system with non-trivial far-field. In a reminiscent analysis, the regularity and integrability properties of its unknowns $(\rho,J)$ corresponding to the mass density $\rho=|\psi|^{2}$ and momentum density $J=\operatorname{Im}(\overline{\psi}\nabla\psi)$ are then captured in terms of Orlicz spaces, see [3] and [39, Chapter 2] as well as [2, 40] for the respective uniform bounds for solutions to the quantum Navier-Stokes equations, a viscous regularization of the QHD system.
@@ 2.3. Smooth approximation
The first statement is proven in [26, Lemma 6] by considering the convolution with a standard mollification kernel and the second statement follows from [25, Proposition $1.$]. In [26, 25], the statements are given for $(\mathbb{E}_{\mathrm{GL}},d_{\mathbb{E}_{\mathrm{GL}}})$ being equivalent to $(\mathbb{E},d_{\mathbb{E}})$ by virtue of Lemma 2.6.
@@ 2.4. Action of the linear propagator on the energy space
The action of the linear Schrödinger group on the space $X^{k}({\mathbf{R}}^{d})+H^{k}({\mathbf{R}}^{d})$ is well-defined, see [26, Lemma 3] and also [27]. While the results in [26, 27] are stated for $(\mathbb{E}_{\mathrm{GL}},d_{\mathbb{E}_{\mathrm{GL}}})$, we state them $(\mathbb{E},d_{\mathbb{E}})$ which by Lemma 2.6 is equivalent.
@@
(2.19) $\|\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}f\|_{X^{k}+H^{k}}\leq C\left(1+|t|\right)^{\frac{1}{2}}\|f\|_{X^{k}+H^{k}},$
@@
The action of the linear Schrödinger group on the space $\mathbb{E}({\mathbf{R}}^{d})$ is described by [26, Proposition 2.3].
@@
Let $d=2,3$. For every $t\in{\mathbf{R}}$, the linear propagator $\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}$ maps $\mathbb{E}({\mathbf{R}}^{d})$ to itself and for every $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{d})$ the map $t\in{\mathbf{R}}\mapsto\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}\in\mathbb{E}({\mathbf{R}}^{d})$ is continuous. Moreover, given $R>0$, $T>0$ there exists $C>0$ such that for every $\psi_{0}^{1},\psi_{0}^{2}\in\mathbb{E}({\mathbf{R}}^{d})$ with $\mathcal{E}(\psi_{0}^{1})\leq R$, $\mathcal{E}(\psi_{0}^{2})\leq R$ one has
@@
Furthermore, given $R>0$, there exists $T(R)>0$ such that, for every $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{d})$ with $\mathcal{E}(\psi_{0})\leq R$, we have
@@ 2.5. Strichartz estimates
for any admissible pair $(q_{1},r_{1})$. Further, in order to discuss the well-posedness theory for (1.1) in the energy space, we also work with the function space $S^{1}(I\times{\mathbf{R}}^{d})$ and $N^{1}(I\times{\mathbf{R}}^{d})$ defined by the norms
@@
Let $T>0$ and $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{d})$, then Lemma 2.14 states that for any admissible pair $(q,r)$, we have
@@
Observe that, by virtue of Lemma 2.11, one has $\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}-\psi_{0}\in C([0,T];H^{1}({\mathbf{R}}^{d}))$ and $\nabla\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}\in C([0,T];L^{2}({\mathbf{R}}^{d}))\cap S^{0}([0,T]\times{\mathbf{R}}^{d}))$.
@@ 2.6. The nonlinearity
(2.30) $\mathcal{N}_{2,\mathrm{bd}}=\mathcal{N}_{2}(\psi)\chi(2\psi),\quad\mathcal{N}_{2,\mathrm{int}}(\psi)=\mathcal{N}_{2}(\psi)(1-\chi(2\psi))$
@@
$|\mathcal{N}_{2,\mathrm{bd}}(\psi)|$ $\leq C(1-\eta(|\psi|),\quad|\mathcal{N}_{2,\mathrm{int}}(\psi)|\leq C|\psi|^{2\alpha+1}(1-\chi(\psi)).$
In the case of vanishing boundary conditions and infinity, the strategy developed in [42], see also [14, Chapter 4], relies on similar pointwise bounds on $\mathcal{N}$. However, here we need to consider additional cut-off functions $\eta$ isolating the behavior close to $1$ in view of the far-field and the related support properties. Note that (2.4) yields that the measure of $\operatorname{supp}(\mathcal{N}_{2}(\psi))$ is bounded by $\mathcal{E}(\psi)$. The quantity $\nabla\mathcal{N}$ can be rigorously defined by means of Nemicki operators, see [42, Appendix A] and also [43, 14]. It reads
@@
For general $\psi_{1},\psi_{2}\in\mathbb{E}({\mathbf{R}}^{d})$ one has $\psi_{1}-\psi_{2}\notin L^{p}({\mathbf{R}}^{d})$ for any $p\geq 1$, unless $\psi_{1},\psi_{2}$ belong to the same connected component of $\mathbb{E}({\mathbf{R}}^{d})$, see Remarks 2.3 and 2.4 for $d=2,3$ respectively. This motivates the following estimates,
@@
$\left|\mathcal{N}_{2,\mathrm{bd}}(\psi_{1})-\mathcal{N}_{2,\mathrm{bd}}(\psi_{2})\right|$ $\leq C\left|\psi_{1}-\psi_{2}\right|,$
$\left|\mathcal{N}_{2,\mathrm{int}}(\psi_{1})-\mathcal{N}_{2,\mathrm{int}}(\psi_{2})\right|$ $\leq C\left(|\psi_{1}|^{2\alpha}+|\psi_{2}|^{2\alpha}\right)\left|\psi_{1}-\psi_{2}\right|.$
@@
(2.36) $G_{1}(\psi)=f(|\psi|^{2})+f^{\prime}(|\psi|^{2})|\psi|^{2},\quad\text{and}\quad G_{2}(\psi)=f^{\prime}(|\psi|^{2})\psi^{2}.$
@@
$G_{i,\mathrm{bd}}(\psi):=G_{i}(\psi)\chi(\psi),\quad\text{and}\quad G_{i,\mathrm{int}}(\psi):=G_{i}(\psi)(1-\chi(\psi)),$
@@
(2.37) $G_{\mathrm{bd}}:=G_{1,\mathrm{bd}}+G_{2,\mathrm{bd}},\quad\text{and}\quad G_{\mathrm{int}}:=G_{1,\mathrm{int}}+G_{2,\mathrm{int}}.$
@@
(2.38) $|G_{\mathrm{bd}}(\psi)|\leq C,\quad\text{and}\quad|G_{\mathrm{int}}(\psi)|\leq C(1+|\psi_{\mathrm{int}}|^{2\alpha})(1-\chi(\psi)).$